OpenStudy (anonymous):

Use synthetic division amd the given factor to complete factor each polynomial function y=x^3 +2x^2-5x-6;(x+1)

OpenStudy (anonymous):

x^2+x-6 r=0

OpenStudy (anonymous):

hope that helped :)

OpenStudy (anonymous):

so you find the coefficent for all of the parts....

OpenStudy (anonymous):

the coefficent for x^3 is 1 since there isn't a number infront of x the coefficent for 2x^2 is 2 the coefficent for -5x is -5 and the coefficent for -6 is -6

OpenStudy (anonymous):

after you find that you solve for (x+1) and you get x=-1

OpenStudy (anonymous):

get it so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oky then you set it up like this

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

after you do that you do this- you bring down 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then you muliply -1 and 1 like this and then subtract like in normal division

OpenStudy (anonymous):

ok I have a quick ? on the second step after I find the coefficient where do you gt the x=-1?

OpenStudy (anonymous):

you solve for x so you subtrat the 1 getting x=-1

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so next you do this

OpenStudy (anonymous):

got it. Thank you so much. Ok i just did this one on my own. tell me if I have it right

OpenStudy (anonymous):

kay show me

OpenStudy (anonymous):

y=x^3-4x^3-4x^2-9x+36; (x+3) my answer to this ons x^2+x-12=r=0

OpenStudy (anonymous):

i got -3x^2+5x-24 r 108

OpenStudy (anonymous):

:-( grrr let me try to see if I come up with your answer to see whre I mess up

OpenStudy (anonymous):

don't forget to simplify

OpenStudy (anonymous):

my coefficients are 1-4-9+36

OpenStudy (anonymous):

ok x^3 - 4x^3 = -3x^3

OpenStudy (anonymous):

omg I'm getting frustrate it.

OpenStudy (anonymous):

lol, just take a deep breathe it will come to you soon, remember if there is a minus sign in front of a number it means it's negative so you could always think of it as 1 + (-4)

OpenStudy (anonymous):

Mariah thanks for help. Im going to bed is late. I'll ask my teacher tomorrow cus I cant think right now. once again thank you so much. i'll log in back tomorrow if I need help again. good night

OpenStudy (anonymous):

gnight btw guy or girl?

OpenStudy (anonymous):

guy

OpenStudy (anonymous):

cool, and what hs do you go to?

OpenStudy (anonymous):

University HS n U

OpenStudy (anonymous):

Kentwood hs, wa where r u?

OpenStudy (anonymous):

Florida

OpenStudy (anonymous):

sad days, i think we would be good friends :) im a girl btw

OpenStudy (anonymous):

why sad dys?

OpenStudy (anonymous):

because i live in washington state... though i do have family in Texas, unless we talk via email i wont be able to talk to you

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

oh ok. thats true. ill log in tomorrow hopefully you gt to help me with my alg issues. GN and nice meeting you

OpenStudy (anonymous):

nice meeting you too >.< hopefully i'll be on too if not horse2503 at yahoo :D

OpenStudy (anonymous):

got btw my name is Giovanni

OpenStudy (anonymous):

nice name :D

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

np... Night :))

OpenStudy (anonymous):

gn :-)